# LUT-LLM: Efficient Large Language Model Inference with Memory-based Computations on FPGAs

### v1

@@ 4.5. Other Design Features
Efficient Attention for both Prefill and Decode Stages. Since LUT-LLM performs the prefill and decode stage on the same design, ensuring high efficiency and compatibility for both stages simultaneously can be challenging. The prefill stage can parallelize computations along the sequence for $Q$, $K$, and $V$, while the decode stage only has a single token $Q$ vector involved. In LUT-LLM, we parallelize computations only along the hidden dimension and sequence dimension for $K$ and $V$. This ensures full utilization of every compute element in both prefill and decode stages.
@@
KV Cache Prefetch and Write-out Orchestrations. During prefill stage, the KV embeddings are streamed to the HBM controller while delivering to the dataflow attention engine. During the decode stage, previously dumped KV cache will be prefetched from HBM when the first linear projection is running. After the KV embeddings for the new token are produced, it will be written to HBM at the end of the KV cache.
5. Experiment
5.1. Experiment Setup
@@
Models, Datasets, and Algorithm. Our overall scheme (following the notation in Figure 1) is: $G=512$, $v=2$, $c_{w}=16$, $c_{a}=64$, and INT8 lookup tables. This scheme is equivalent to W2A3. Based on Qwen 3 1.7B model (yang2025qwen3), we first finetune for activation quantization. We applied per-tensor zero-point quantization that converts FP32 lookup tables to unsigned INT8. This quantizes the FP32 matrix $X$ into quantized matrix $X_{INT8}$ by
(13) $\begin{split}s=(\max(X)-\min(X))/256,z=-\min(X)/s\\ X_{INT8}=\min(256,\max(0,sX+z))\end{split}$
@@
We follow the two-stage training strategy for centroids and weights in LUT-DLA (li2025lut) with reconstruction loss ratio as 0.1. Additionally, we made algorithmic optimizations for the forward and backward passes. In the forward pass, we construct a new lookup table gather reduce kernel, which cuts memory usage with runtime lookup table reconstructions and accelerates the training process by fusing table lookups and accumulations. In the backward pass, we fuse the kernels that compute the gradient of weights and centroids. These updates allow us to train models beyond 1B parameters. Quantization-aware training is supported with the straight-through estimator (STE) (yin2019understanding) with adjustable gradients. The training for activation vector quantization costs more GPU resources than conventional language model training since we need to tune all lookup tables in full precision and introduce extra computation burdens for quantization, dequantization, and reconstruction. In our setting, an eight-GPU cluster will take 2 weeks to first pretrain Qwen 3 1.7B on FineWeb (penedo2024fineweb) for 512-token sequences and then finetune on WikiQA (yang2015wikiqa) with 3 epochs, with 75% memory utilization. Scaling up to larger models and longer sequences requires more GPU memory and time, or further optimized kernels.
@@
After training for activation quantization, we reconstruct the weights from the trained lookup tables, then apply GPTVQ (van2024gptvq). Then, we then pre-compute the new lookup tables by computing the dot products between the vectors in activation codebooks and weight codebooks.
@@
For evaluation, we utilize the GLUE benchmark (wang2018glue), SQuAD v2 (rajpurkar2016squad), and MMLU-Pro (wang2024mmlu) for question-answering dataset. Benchmarks evaluate performance based on accuracy and F1 score. We compare with SmoothQuant (xiao2023smoothquant) and GPTQ (frantar2022gptq) as the representatives of scalar quantization schemes.
Metrics. To compare the accelerator performance between LUT-LLM, GPUs, and other accelerators, we employ latency in miliseconds and energy efficiency in tokens per Joule as the metrics. We measure both end-to-end latency and decoding latency for better understanding of scaling to longer sequence. Energy efficiency measures the number of tokens generated per energy cost.
@@
Target Devices and Tools. LUT-LLM is prototyped on AMD Alveo V80 FPGA (amd2024alveov80). To design the circuit, we use Xilinx Vitis HLS 2024.2 with TAPA framework (guo2023tapa) and utilize RapidStream (guo2022rapidstream; guo2021autobridge) for coarse-grain floorplanning to reduce the routing congestion and improve timing with register insertions. The design is implemented with Vivado 2024.2 with our customized block design to integrate the HLS-based RTL block and generate the bitstream.
Figure 11. End-to-end and decode latency of LUT-LLM and GPUs with BF16 and INT8 precisions. Context window is within 512 tokens due to the limitation of training algorithm and system for longer context.
Figure 12. Energy efficiency (tokens per Joule) of LUT-LLM and GPUs with BF16 and INT8 precision.
GPU Benchmarking. We select two GPUs with a similar process as V80 FPGA: AMD Instinct MI210 (amd2025mi210) and NVIDIA A100 (nvidia2021a100). They have different hardware configurations and compiler optimizations, exhibiting various behaviors in language model inference performance. Table 2 compares V80 with the GPUs. The peak TOPs for V80 is scaled based on the achieved frequency. We use vLLM (kwon2023efficient) as the serving backend for benchmarking. We enable FlashAttention (dao2023flashattention) and FlashDecoding (hong2023flashdecoding++) to maximize the GPU performance with KV cache optimizations. Besides the model in the original precision, we also deploy the INT8 model quantized with GPTQ (frantar2022gptq), which is the lowest precision available for Qwen 3. A100 has GPTQ Marlin kernel (frantar2025marlin) integrated for efficient INT8 inference. MI210 does not support this optimization. We write a python script using pyNVML (nvidia-ml-py) and pyrsmi to monitor the GPU power consumption.
Table 2. Hardware Configuration of FPGA and GPUs
@@
Table 3. Evaluation on GLUE and SQuAD v2 datasets for Qwen 3 1.7B with different quantization schemes.
@@
MNLI MRPC QNLI QQP RTE SST-2 STS-B SQuAD v2
FP16 Baseline 87.56 86.52 92.89 91.18 80.87 93.69 87.66 72.83
RTN INT8 86.67 80.15 87.99 89.26 70.40 87.39 77.27 62.00
SmoothQuant (xiao2023smoothquant) W8A8 86.99 85.29 91.72 89.75 79.06 91.17 85.93 71.26
LUT-LLM:
Act. Quant. 87.02 84.07 91.91 90.69 78.34 91.17 85.56 70.29
+ INT8 LUT 86.93 83.82 91.71 90.82 76.89 90.71 84.87 69.82
+ Weight Quant. 86.91 82.84 90.35 89.54 76.53 90.6 85.07 69.73
Table 4. Evaluation on MMLU-Pro benchmark for Qwen 3 1.7B with different quantization schemes.
Model MMLU-Pro (Avg)
FP16 Baseline 33.1
W8A16 GPTQ 32.1
W8A8 SmoothQuant 31.7
LUT-LLM 30.8
@@
SoTA FPGA Accelerator Baselines. We select two representative SoTA works, Allo (chen2024allo), FlightLLM (zeng2024flightllm), and InTAR (he2025intar). They are arithmetic-based and implemented on the Alveo U280 (amd2023alveou280). Allo and InTAR applies W4A8 scheme and FlightLLM utilizes 3.5-bit sparse weights. FlightLLM employs aggressive sparsification and quantization on both weights and attention. This fundamentally alters the computation pattern and data loading size. Since the mapping flow and quantization scripts are not open-sourced, we use the performance model in FlightLLM’s artifact to estimate the latency. Since these accelerators are implemented on U280 and do not natively support the latest Qwen 3 1.7B model, we construct a simulator based on their designs, with less than a 2% gap from data in the original papers.
5.2. Algorithm Performance
With the quantization scheme of LUT-LLM, the target model maintains a competitive performance compared with the FP16 baseline and consistently outperforms standard round-to-nearest (RTN) quantization to INT8 for activations. Table 3 summarizes the model quality on GLUE and SQuAD v2 under different quantization schemes. LUT-LLM incurs only a $2.7\%$ performance drop relative to the FP16 baseline with all techniques applied. Compared with RTN INT8 quantization, LUT-LLM with INT8 lookup tables achieves $5.5\%$ higher performance in geomean. Compare with SmoothQuant, LUT-LLM has only $1.2\%$ drop in geomean accuracy while applying more aggressive quantization scheme (LUT-LLM is equivalent to W2A3). We also benchmarked LUT-LLM on harder benchmarks, such as MMLU-Pro, achieving only $0.9\%$ accuracy drop than SmoothQuant. These results highlight the effectiveness of the quantization scheme and tuning methods used in LUT-LLM in preserving model quality while enabling aggressive quantization for efficient deployment on FPGAs.
5.3. Accelerator Performance
@@
Figure 13. LUT-LLM implementation layout on AMD V80 FPGA, resource utilization, and clock frequency.
5.3.1. Resource Utilization
@@
Figure 13 presents the floorplan of LUT-LLM on the AMD V80 FPGA. Since LUT-LLM primarily targets improving decoding and short-context prefilling efficiency, both of which are bounded by linear layers, a substantial portion of resources is allocated to the LUTLinear engine. The design attains a clock frequency of 227 MHz. Although the utilization of all resource types remains at or below 50%, the CLB utilization reaches 81%. This imbalance, caused by resource heterogeneity, imposes significant challenges in placement and routing. LUT-LLM employs 32 HBM channels for reading lookup tables and weight indices, along with 4 channels for input data and 4 channels for outputs and KV cache. Each channel on the V80 provides a 256-bit interface. The HBM clock is aligned with the kernel frequency to sustain a balanced read and write throughput. Consequently, the effective bandwidth utilized for loading lookup tables is $227\times 256\times 32/8/1024=227~\text{GB/s}$.
5.3.2. Comparison with GPUs
We compare the end-to-end latency of LUT-LLM with GPUs running Qwen 3 1.7B in both BF16 and INT8 precisions. Figure 11 shows the latency breakdown of the prefill and decode stages for each configuration. Due to training memory constraints on GPUs, the vector-quantized model can currently process up to 512 tokens, and we therefore report results within this context window. With INT8 quantized lookup tables, LUT-LLM has a $1.66\times$ lower geomean latency than MI210, despite utilizing $7\times$ less memory bandwidth. For the A100, the FlashAttention and FlashDecoding algorithms deliver substantial speedups in both prefill and decode stages, and the GPTQ Marlin kernel further accelerates INT8 inference on tensor cores. As a result, LUT-LLM is slower. However, considering that LUT-LLM consumes $8.8\times$ less bandwidth than the A100 and decoding is fundamentally memory-bound, this modest latency gap is still notable. A detailed latency analysis reveals that LUT-LLM decoding saturates 92% of the available bandwidth, approaching the hardware limit. Furthermore, we show the performance of LUT-LLM with INT4 lookup tables without altering computational resources. In this case, the latency aligns with the A100 in BF16 and INT8 as the output length increases.
@@ 5.3.2. Comparison with GPUs
We then compare the energy efficiency, as shown in Figure 12. While all GPUs benefit from INT8 quantization due to reduced off-chip memory traffic, LUT-LLM maintains a substantial advantage, achieving $4.1\times$ higher geomean energy efficiency than MI210. On the A100, the GPTQ Marlin kernel not only reduces off-chip memory access but enhances INT8 computation efficiency, narrowing the gap relative to LUT-LLM. Nevertheless, LUT-LLM remains $1.72\times$ more energy efficient in geomean. Moreover, the efficiency gap between LUT-LLM and GPUs widens as output length increases, underscoring the scalability of LUT-LLM for long output sequence.
5.3.3. Comparison with SoTA FPGA Accelerators
Figure 14. Inference latency comparison of LUT-LLM, InTAR, Allo, and FlightLLM.
Table 5. Comparison of energy efficiency across accelerators.
Accelerator Geomean Energy Efficiency (tokens/J)
FlightLLM 1.54
InTAR 1.67
Allo 0.69
LUT-LLM 2.29
@@ 5.3.3. Comparison with SoTA FPGA Accelerators
Figure 14 compares the end-to-end latency of LUT-LLM with InTAR, Allo, and FlightLLM. Overall, LUT-LLM is $5.6\times$ and $1.9\times$ faster than Allo and InTAR, respectively. The performance gain arises from improvements in both the prefill and decode stages. In the prefill stage, LUT-LLM leverages memory-based table lookups as computation, enabling higher throughput than Allo and InTAR, which are arithmetic-based accelerators for quantized models. As the input length increases, the latency gap between LUT-LLM and InTAR during the prefill stage narrows, driven by the costly FP32 attention computations and the reduced resource allocation for attention needed to sustain high throughput in the linear layers. In the decode stage, LUT-LLM performs computations more efficiently, fully utilizing the available memory bandwidth. Moreover, Allo and InTAR incur additional overhead from dequantization, which requires extra data loading and costly floating-point computations. Consequently, even when weights are loaded in INT4 precision, the decoding speed of Allo and InTAR is slower than that of LUT-LLM, which streams INT8 data from HBM. Compared with FlightLLM, LUT-LLM has similar end-to-end latency and $1.1\times$ faster decoding speed, even with more data loaded from HBM. This indicates a better scaling of LUT-LLM with respect to the output length.
Table 5 illustrates the energy efficiency of each FPGA accelerators, where the energy profile of FlightLLM, InTAR, and Allo are from the U280 FPGA. LUT-LLM is $1.49\times$, $1.37\times$, and $3.32\times$ more energy efficient than FlightLLM, InTAR, and Allo, respectively, demonstrating promising energy saving of memory-based computes on FPGA-based LLM acceleration.
5.3.4. Discussion: Model Scaling and Adaptation
Although the current training algorithm limits the model size, LUT-LLM can deploy larger models with the same quantization scheme with simple changes of loop bounds in our design. Overall, LUT-LLM is $2.16\times$ more energy efficient than the A100. LUT-LLM is also parametrized for different quantization schemes and model architecture for fast accelerator generation, but the optimization requires manual design. We leave the automation as the future work.
6. Conclusion and Outlook
This work presents LUT-LLM, the first FPGA accelerator for LLM inference based on memory-driven computation. We develop a performance model for vector-quantized language models on FPGAs and demonstrate that co-quantizing activations and weights is essential to surpass conventional scalar quantization, realized through efficient 2D table lookups. LUT-LLM achieves $1.72\times$$4.1\times$ higher energy efficiency than GPUs with state-of-the-art optimization algorithms and low-precision tensor cores. Future work will further enhance efficiency through emerging algorithmic advances, such as low-bit attention (liu2024spinquant) and low-bit centroid search. In addition, scaling memory bandwidth to GPU levels can deliver up to a $3.19\times$ speedup over the A100, suggesting promising directions in multi-FPGA systems and next-generation HBM technology on FPGAs for LUT-LLM.
7. Appendix: Performance Derivation

### v2

@@ IV-E Other Design Features
Efficient Attention for both Prefill and Decode Stages. Since LUT-LLM performs the prefill and decode stage on the same design, ensuring high efficiency for both stages simultaneously can be challenging. The prefill stage can parallelize computations along the sequence for $Q$, $K$, and $V$, while the decode stage only has a single token $Q$ vector involved. In LUT-LLM, we parallelize computations only along the hidden dimension and sequence dimension for $K$ and $V$. This ensures full utilization of every compute element in both stages.
@@
KV Cache Prefetch and Write-out Orchestrations. During the prefill stage, the KV cache is streamed to the HBM controller while delivering to the dataflow attention engine. During the decode stage, previously dumped KV cache is prefetched from HBM when the first linear projection is running. The KV embeddings of the new token will then be written to HBM at the end of the KV cache.
V Experiment
V-A Experiment Setup
@@
Models, Datasets, and Algorithm. Our overall scheme (following the notation in Table I) is: $G=512$, $v=2$, $c_{w}=16$, $c_{a}=64$, and INT8 quantized lookup tables. Based on Qwen 3 1.7B model [60], we first calibrate for activation quantization. We applied per-tensor zero-point quantization [26] on the lookup tables, which quantizes the FP32 lookup tables $X$ into quantized matrix $X_{INT8}$ by
$\begin{split}s=(\max(X)-\min(X))/256,z=-\min(X)/s\\ X_{INT8}=\min(256,\max(0,sX+z))\end{split}$ (10)
@@
We optimize the two-stage training scheme of LUT-DLA [30]. In the first stage, a fine-grained, layer-wise initialization of K-means is applied to obtain the activation centroids. This improves the overall training stability. Additionally, we design algorithmic optimizations for the forward and backward passes. In the forward pass, we construct a new lookup table gathering reduce kernel, which cuts memory usage with runtime lookup table reconstructions and accelerates the training process by fusing table lookups and accumulations. In the backward pass, we fuse the kernels that compute the gradient of weights and centroids. These updates allow us to train models beyond 1B parameters. Quantization-aware training is supported with the Straight-Through Estimator (STE) [62] with adjustable gradients. With these optimizations, the total training time is reduced from over 1,000 A100 GPU hours to approximately 10 A100 GPU hours, representing an efficiency improvement over LUT-DLA.
@@
After training for activation quantization, we reconstruct the weights from the trained lookup tables, then apply GPTVQ [50]. Then, we pre-compute the new lookup tables by computing the dot products between the vectors in activation codebooks and weight codebooks.
@@
For evaluation, we utilize the GLUE benchmark [52], SQuAD v2 [46], and MMLU-Pro [56] for question-answering dataset. Benchmarks are evaluated based on accuracy and F1 score. We compare with SmoothQuant [59] and SpinQuant [36] in W8A8 to represent scalar quantization schemes.
Metrics. We employ latency (end-to-end and decoding) and energy efficiency in tokens per Joule as metrics. We convert latency to relative speedup for better visualization.
@@
Target Devices and Tools. LUT-LLM is prototyped on AMD Alveo V80 FPGA [5]. To design the circuit, we use Xilinx Vitis HLS 2024.2 with TAPA framework [19] and utilize RapidStream [21, 20] for coarse-grain floorplanning to reduce routing congestion and improve timing. The design is implemented with Vivado 2024.2 with our customized block design to integrate the HLS-based RTL block and generate the bitstream. Resource utilizations is in Table LABEL:tab:resource.
GPU Benchmarking. We select AMD Instinct MI210 [4] and NVIDIA A100 [40] as representatives. Table II compares V80 with the GPUs. The peak TOPs for V80 is scaled based on the achieved frequency. We use Huggingface Transformers for a naive GPU baseline and vLLM [28] with GPTQ [15] INT8 quantization as an optimized baseline for benchmarking. A100 has GPTQ Marlin kernels [16] integrated for efficient INT8 and INT4 inference. MI210 does not support this optimization. We utilize pyNVML [41] and pyrsmi to monitor GPU power.
TABLE II: Hardware Configuration of FPGA and GPUs
@@
TABLE III: Evaluation on GLUE, SQuAD v2, and MMLU-Pro datasets for Qwen 3 1.7B with different quantization schemes.
@@
MNLI MRPC QNLI QQP RTE SST-2 SQuADv2 MMLU-Pro
FP16 Baseline 87.6 86.5 92.9 91.2 80.9 93.7 72.8 33.1
RTN INT8 86.7 80.2 88.0 89.3 70.4 87.4 62.0 23.6
SmoothQuant [59] 87.0 85.3 91.7 89.6 79.1 91.2 71.3 31.7
SpinQuant [36] 87.3 83.3 91.8 89.5 80.2 91.5 72.0 28.0
LUT-LLM
+ Act. Quant. 87.0 84.1 91.9 90.7 78.3 91.2 70.3 31.8
+ INT8 LUT 86.9 83.8 91.7 90.8 76.9 90.7 69.8 31.3
+ Weight Quant. 86.9 82.8 90.4 89.5 76.5 90.6 69.7 30.8
@@
SoTA FPGA Accelerator Baselines. We select three representative SoTA works, Allo [10], FlightLLM [63], and InTAR [22]. They are arithmetic-based accelerators. Allo and InTAR apply the W4A8 scheme and FlightLLM utilizes 3.5-bit weights. FlightLLM employs aggressive sparsification and quantization on both weights and attention. This fundamentally alters the computation pattern and data loading size. Since these accelerators do not natively support the latest Qwen 3 1.7B model, we construct a simulator for Allo and InTAR and modify the simulator in FlightLLM’s artifact to get their latency on V80, with less than a 2% gap from data in the original papers.
@@
Figure 10: LUT-LLM implementation layout on AMD V80 FPGA, resource utilization, and clock frequency.
V-B Algorithm Performance
With the quantization scheme of LUT-LLM, the target model maintains a competitive performance compared with the FP16 baseline and consistently outperforms standard round-to-nearest (RTN) quantization to INT8 for activations. Table III summarizes the model quality under different quantization schemes. Compared with SoTA scalar quantization technique, SpinQuant [36], LUT-LLM has only $1.8\%$ drop in geomean accuracy. These results highlight the effectiveness of the quantization scheme and tuning methods used in LUT-LLM in preserving model quality and efficient deployment on FPGAs.
V-C Accelerator Performance
Figure 11: End-to-end and decode speedup of LUT-LLM and GPUs (both in 7nm process node) with BF16, INT8, and INT4 precisions.
Figure 12: Energy efficiency (tokens per Joule) of LUT-LLM and GPUs with BF16, INT8, and INT4 precision.
V-C1 Resource Utilization
@@
Figure 10 presents the floorplan of LUT-LLM on the AMD V80 FPGA. The design attains a clock frequency of 250 MHz. Although the utilization of all resource types remains at or below 50%, the CLB utilization reaches 80%. LUT-LLM employs 32 HBM channels for reading lookup tables and weight indices, along with 4 channels for input data and 4 channels for outputs and KV cache. Each channel on the V80 provides a 256-bit interface. The HBM clock is aligned with the kernel frequency to sustain a balanced read and write throughput. Consequently, the effective bandwidth utilized for loading lookup tables is $250\times 256\times 32/8/1024=250~\text{GB/s}$.
V-C2 Comparison with GPUs
Figure 11 reports the end-to-end latency and decoding speedup across all configurations. Leveraging quantized lookup tables, LUT-LLM achieves a $3.29\times$ lower geometric-mean latency than the MI210 with INT8 quantization. On the A100, LUT-LLM delivers a $1.46\times$ speedup over the FP16 baseline and $1.21\times$ over the INT8 baseline. Notably, even compared to A100 with INT4 quantization, which reduces model storage by half relative to LUT-LLM, our approach still attains a $1.10\times$ decoding speedup.
The limited speedup of INT4 on GPUs compared to INT8 can be attributed to two factors. First, vLLM employs the Marlin kernel for INT4 execution, which performs on-the-fly dequantization from INT4 to FP16 and subsequently utilizes FP16 tensor core instructions for multiply-accumulate operations. This dequantization overhead and reliance on higher-precision compute reduce the achievable throughput compared to native low-precision execution. Second, effective HBM bandwidth utilization on GPUs degrades for smaller models at lower precision. We observed that with a 1.7B model in INT4, the A100 attains only $0.6\times$ of the bandwidth utilization than with INT8. We acknowledge that this diminished scaling may not hold for larger models [53]: when extrapolated to Qwen 3 32B, LUT-LLM achieves $1.2\times$ speedup over A100 with INT8, but only $0.8\times$ when deploying the INT4 model on A100.
@@ V-C2 Comparison with GPUs
We then compare the energy efficiency, as shown in Figure 12. While all GPUs benefit from quantization due to reduced off-chip memory traffic and specialized low-bit tensor core, LUT-LLM maintains a substantial advantage, achieving $6.6\times$ higher geomean energy efficiency than MI210. For A100, LUT-LLM remains $3.05\sim 5.94\times$ more energy efficient in geomean. Moreover, the efficiency gap between LUT-LLM and GPUs widens as output length increases, underscoring the scalability of LUT-LLM for long output sequences.
V-C3 Comparison with SoTA FPGA Accelerators
Figure 13: Speedup of LUT-LLM over InTAR, Allo, and FlightLLM.
@@ V-C3 Comparison with SoTA FPGA Accelerators
Figure 13 compares the latency of LUT-LLM with InTAR, Allo, and FlightLLM. Overall, LUT-LLM is $5.6\times$ and $1.9\times$ faster than Allo and InTAR. The performance gain arises from improvements in both the prefill and decode stages. In the prefill stage, LUT-LLM leverages on-chip memory access as computation for higher peak TOPs than Allo and InTAR. In the decode stage, LUT-LLM performs computations more efficiently, fully utilizing the available memory bandwidth. Moreover, Allo and InTAR incur additional overhead from dequantization, which requires extra data loading and costly floating-point computations. Compared with FlightLLM, LUT-LLM is $1.6\times$ faster than FlightLLM end-to-end, with larger speedup as sequence length grows.
VI Conclusion and Outlooks
This work presents LUT-LLM, the first FPGA accelerator for 1B+ LLM inference with memory-based computation. We develop a performance model and demonstrate that co-quantization is essential for speedup. LUT-LLM achieves $1.10\sim 3.29\times$ higher speed than GPUs. LUT-LLM can be mapped to ASICs by replacing FPGA-specific IPs. It is also promising for future memory-stacked designs [12].
Acknowledgement
This work was supported in part by PRISM, one of the seven centers in the JUMP 2.0 program sponsored by SRC and DARPA. It is also supported by CDSC industrial partners and the AMD  [^2] J. Cong has a financial interest in AMD HACC Program.
VII Appendix: Performance Derivation